OpenStudy (anonymous):

Solve. 3(3a-2)=5/8(8- 32 a )

OpenStudy (anonymous):

3(3a-2) = 5/8(8-32a) 3(3a) + 3(-2) = (5/8)(8) + (5/8)(-32a) 9a - 6 = (5x8/8) + ( -32a x 5 / 8 ) 9a - 6 = 5 - 20a 9a + 20a = 5 + 6 29a = 11 a = 11/29